# The Diophantine problem in isotropic reductive groups

### v1

@@ 4 Some normalizers and centralizers
$\operatorname{C}_{G}(Z_{\alpha})\leq\begin{cases}G^{0}_{\{\alpha,2\alpha\}}&\text{ if $\Phi$ is of type $\mathsf{BC}_{\ell}$, $\ell\geq 2$, $\alpha$ is ultrashort};\\ G^{0}_{\{2\mathrm{e}_{i},\mathrm{e}_{i}+\mathrm{e}_{j},2\mathrm{e}_{j}\}}&\text{ if $\Phi$ is of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$, $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ is short};\\ G^{0}_{\alpha}&\text{ otherwise}.\end{cases}$
It remains to show that $\operatorname{C}_{G}(Z_{\alpha})\cap L=\mathrm{C}(G)$ or $\mathrm{C}^{\mathrm{us}}(G)$. If $\alpha$ is not ultrashort we are done by lemma 5 applied to $f(\beta)=2\frac{\alpha\cdot\beta}{\alpha\cdot\alpha}$. Otherwise note that $\Gamma_{\alpha}$ is a parabolic subset of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$. Applying lemma 5 to $G^{0}_{\{-\beta,\beta\}}$ for various $\beta\in\Gamma_{\alpha}$ with $f(\gamma)=2\frac{\gamma\cdot\beta}{\beta\cdot\beta}$ (more precisely, to its simple subgroup containing $G_{\beta}$ after a splitting étale extension) we see that $\operatorname{C}_{L}(t_{\beta}(X_{\beta}))$ centralizes both $G_{\beta}$ and $G_{-\beta}$, so $\operatorname{C}_{G}(Z_{\alpha})\cap L=\operatorname{C}^{\mathrm{us}}(G)$. ∎
@@
$\operatorname{C}_{G}\bigl{(}\bigcup_{\alpha\in\Phi}t_{\alpha}(X_{\alpha})\bigr{)}=\operatorname{C}(G),\quad\{g\in G\mid\forall\alpha\in\Phi\enskip{{}^{g}\!{t_{\alpha}(X_{\alpha})}}\subseteq G_{\alpha}\}=L.$
@@
The first claim follows from lemma 2 applied to opposite long roots and lemma 2. The second one is a corollary of theorem 1 applied to all long roots and lemma 2. ∎
@@ 5 Main results
Let $G$ be an isotropic reductive group scheme over $K$ such that the rank of $\Phi$ is at least $2$. Then the sets $L(K)$, $\operatorname{C}(G)(K)$, and all $\operatorname{C}(G)(K)\,G_{\alpha}(K)$ are Diophantine in the group $G(K)$. Moreover, the root subgroups $G_{\alpha}(K)$ are Diophantine in the group $G(K)$ unless the type is $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{BC}_{2}$, or $\alpha$ is short and the type is $\mathsf{G}_{2}$.
@@
By theorem 2 the following groups are Diophantine.
  • $\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\alpha}(K)$ for ultrashort $\alpha$ and $\Phi$ of type $\mathsf{BC}_{\ell}$;
  • $\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{2\mathrm{e}_{j}}(K)$ for $\Phi$ of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$;
  • $\mathrm{C}(G)(K)\,G_{\alpha}(K)$ for $\alpha$ long or $\alpha$ short and $\Phi$ of type $\mathsf{B}_{\ell}$, $\ell\geq 3$, $\mathsf{F}_{4}$, or $\mathsf{G}_{2}$.
Intersections of Diophantine groups, their products and their commutators with single elements are also Diophantine.
@@
$\operatorname{C}(G)(K)=\operatorname{C}(G)(K)\,G_{\alpha}(K)\cap\operatorname{C}(G)(K)\,G_{-\alpha}(K)$
@@
$\{g\in G(K)\mid{{}^{g}\!{t_{\alpha}(X)}}\subseteq\operatorname{C}(G)(K)\,G_{\alpha}(K)\}=G^{0}_{\{\beta\mid\angle(\alpha,\beta)\leq\pi/2\}}(K)$
@@
$G_{\alpha}(K)=\bigl{[}\operatorname{C}(G)(K)\,G_{\beta}(K),t_{\alpha-\beta}(e)\bigr{]}$
is Diophantine, where $\beta$ is long, $\angle(\alpha,\beta)=\frac{\pi}{3}$, and $t_{\alpha-\beta}(e)\in G_{\alpha-\beta}$ is a factor of a Weyl element.
@@
$G_{\alpha}(K)=\operatorname{C}(G)(K)\,G_{\alpha}(K)\cap\bigl{[}\operatorname{C}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$
is Diophantine, where $t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)$ is a factor of a Weyl element. It follows that the same holds for the type $\mathsf{F}_{4}$.
@@
$G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\bigl{[}\bigl{[}\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{j}}(K)\,G_{\mathrm{e}_{j}+\mathrm{e}_{k}}(K)\,G_{2\mathrm{e}_{k}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]},t_{\mathrm{e}_{j}-\mathrm{e}_{k}}(e^{\prime})\bigr{]},$
where constants are factors of Weyl elements. Next, the group
$G_{2\mathrm{e}_{i}}(K)=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\cap\bigl{[}\operatorname{C}(G)\,G_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)]\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$
is Diophantine. If $\Phi$ is of type $\mathsf{BC}_{\ell}$, then
$G_{\mathrm{e}_{i}}(K)=\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\mathrm{e}_{i}}(K)\cap\bigl{[}\operatorname{C}^{\mathrm{us}}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(e)\bigr{]}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{\mathrm{e}_{i}-\mathrm{e}_{j}}(K)$
is also Diophantine.
Finally, consider the exceptional cases $\mathsf{C}_{2}$ and $\mathsf{BC}_{2}$. The group $\operatorname{C}(G)(K)\,G_{\mathrm{e}_{i}}(K)$ is Diophantine (in the case of $\mathsf{BC}_{2}$) since it contains $g\in\operatorname{C}^{\mathrm{us}}(K)\,G_{\mathrm{e}_{i}}(K)$ if and only if
$\bigl{[}g,t_{\mathrm{e}_{i}}(X_{0})\bigr{]}\subseteq\operatorname{C}(G)(K)\,G_{2\mathrm{e}_{i}}(K)$
for some generating set $X_{0}\subseteq\mathfrak{g}_{\mathrm{e}_{i}}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\mathrm{e}_{i}}$. We have
$\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,\prod_{x\in X}\bigl{[}\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}(K)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,G_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr{]}$
for a finite generating set $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ by lemma 3, so $\operatorname{C}(G)\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$ is also Diophantine. ∎
@@
Let $G$ be an isotropic reductive group scheme over $K$ with the root system $\Phi$ of rank at least $2$. Then the ring $K$ is e-interpretable in $G(K)$.
@@
We assume that the type of $\Phi$ is neither $\mathsf{G}_{2}$ nor $\mathsf{BC}_{\ell}$ since these cases easily reduce to $\mathsf{A}_{2}$ and $\mathsf{C}_{\ell}$ respectively. Fix some finite $K$-module generating sets $X_{\alpha}\subseteq\mathfrak{g}_{\alpha}$ for all roots $\alpha$ and let $Y_{\alpha}=X_{\alpha}\cup\bigcup_{\beta,\alpha-\beta\in\Phi}[X_{\beta},X_{\alpha-\beta}]$.
@@
$(k_{\alpha}\colon Y_{\alpha}\to\operatorname{C}(G)(K)\,G_{\alpha}(K)/\operatorname{C}(G)(K))_{\alpha\in\Phi}$
@@
$[k_{\alpha}(y_{\alpha}),t_{\beta}(y_{\beta})]\equiv[t_{\alpha}(y_{\alpha}),k_{\beta}(y_{\beta})]\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
@@
$[k_{\alpha}(x_{\alpha}),t_{\beta}(x_{\beta})]\equiv k_{\alpha+\beta}([x_{\alpha},x_{\beta}])\pmod{\operatorname{C}(G)(K)\,\prod_{\begin{subarray}{c}i\alpha+j\beta\in\Phi\\ i,j\geq 1\\ (i,j)\neq(1,1)\end{subarray}}G_{i\alpha+j\beta}}$
@@
for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$. Here $\alpha$ and $\beta$ are all roots such that $\alpha+\beta$ is also a root. Clearly, $\widetilde{K}$ is the factor-set of a Diophantine set by a Diophantine equivalence relation.
@@
Every $k\in K$ determines a corresponding element $\widetilde{k}\in\widetilde{K}$, $\widetilde{k}(y_{\alpha})=t_{\alpha}(ky_{\alpha})$. Conversely, consider elements $u_{y_{\alpha}}\in\mathfrak{g}_{\alpha}$ for $y_{\alpha}\in Y_{\alpha}$ such that $k_{\alpha}(y_{\alpha})\in t_{\alpha}(y_{x_{\alpha}})\,\operatorname{C}(G)$. For any root $\alpha$ there is long $\beta\in\Phi$ such that $\alpha+\beta\in\Phi$, so $[u_{y_{\alpha}},y_{\beta}]=[y_{\alpha},u_{y_{\beta}}]$. Recall that the Lie bracket $[{-},{=}]\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate (by lemma 3 if $\alpha$ is short). For any linear relation $\sum_{y_{\alpha}\in Y_{\alpha}}y_{\alpha}a_{\alpha}=0$ the equation implies that $\sum_{y_{\alpha}\in Y_{\alpha}}u_{y_{\alpha}}a_{\alpha}=0$, so there is a unique linear map $U_{\alpha}\colon\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}$ such that $u_{y_{\alpha}}=U_{\alpha}(y_{\alpha})$.
@@
The defining congruences imply that $[U_{\alpha}(x),y]=U_{\alpha+\beta}([x,y])=[x,U_{\beta}(y)]$. We claim that such maps $U_{\alpha}$ are necessarily scalar. It suffices to consider $\Phi$ of rank $2$ and to check the claim for only one root $\alpha$. This is clear if one of $\mathfrak{g}_{\alpha}$ is one-dimensional.
  • In the linear-like case (${{}^{1}\!{\mathsf{E}_{6,2}^{28}}}$ and ${{}^{1}\!{\mathsf{A}_{3d-1,2}^{(d)}}}$) we have
    $U_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=U_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,U_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$

### v2

@@ 4 Some normalizers and centralizers
[…]
It remains to show that $\operatorname{C}_{L}(Z_{\alpha})=\mathrm{C}(G)$ or $\mathrm{C}^{\mathrm{us}}(G)$. If $\alpha$ is neither ultrashort nor short in $\Phi=\mathsf{G}_{2}$, we are done by lemma 5 applied to $f(\beta)=2\frac{\alpha\cdot\beta}{\alpha\cdot\alpha}$. In the case of $\mathsf{G}_{2}$ first apply the same argument to $G_{\{-\alpha,\alpha\}}$ to get $\operatorname{C}_{L}(Z_{\alpha})\leq\operatorname{C}(L)$ (using $L,L]\leq G_{\{-\alpha,\alpha\}}$), and then note that $\operatorname{C}_{\operatorname{C}(L)}(Z_{\alpha})=\operatorname{C}(G)$ because the preimage of $\Gamma_{\alpha}$ in $\widetilde{\Phi}$ spans the whole root lattice e.g. by considering root diagrams [22].
It remains to consider $\Phi=\mathsf{BC}_{\ell}$ and ultrashort $\alpha$. Note that $\Gamma_{\alpha}$ is a parabolic subset of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$. Applying lemma 5 to $G_{\mathsf{C}_{\ell}}$ with $f(\beta)=2\frac{\beta\cdot\alpha}{\alpha\cdot\alpha}$ we see that $\operatorname{C}_{L}(Z_{\alpha})$ centralizes all $U_{\beta}$ with $\beta\in\mathsf{C}_{\ell}$ (note that $G_{\mathsf{C}_{\ell}}\leq G^{0}_{\mathsf{C}_{\ell}}$ is a simple factor e.g. by [26, lemma 1] and all other factors centralize $G_{\mathsf{C}_{\ell}}\supset Z_{\alpha}$), so $\operatorname{C}_{L}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)$. ∎
@@
$\operatorname{C}_{G}\bigl(\bigcup_{\alpha\in\Phi}t_{\alpha}(X_{\alpha})\bigr)=\operatorname{C}(G),\quad\{g\in G\mid\forall\alpha\in\Phi\enskip{{}^{g}\!{t_{\alpha}(X_{\alpha})}}\subseteq U_{\alpha}\}=L.$
@@
The first claim follows from theorem 2 applied to opposite long roots and lemma 2. The second one is a corollary of theorem 1 applied to all long roots and lemma 2. ∎
@@ 5 Main results
Recall that $\mathrm{E}_{G}(K)=\bigl\langle U_{\alpha}(K)\mid\alpha\in\Phi\bigr\rangle$ is the elementary subgroup [20, 26]. If $\mathrm{E}_{G}(K)\leq H\leq G(K)$ is an intermediate subgroup, then $\operatorname{C}(G(K))\cap H=\operatorname{C}(H)=\operatorname{C}_{H}(\mathrm{E}_{G}(K))$ by theorem 3.
@@
Let $G$ be an isotropic reductive group scheme over $K$ such that the rank of $\Phi$ is at least $2$. Then the sets $L(K)$, $\operatorname{C}(G)(K)$, and all $\operatorname{C}(G)(K)\,U_{\alpha}(K)$ are Diophantine in the group $G(K)$. Moreover, the root subgroups $U_{\alpha}(K)$ are Diophantine in the group $G(K)$ unless the type is $\mathsf{B}_{2}=\mathsf{C}_{2}$, $\mathsf{BC}_{2}$, or $\alpha$ is short and the type is $\mathsf{G}_{2}$. If $\mathrm{E}_{G}(K)\leq H\leq G(K)$ is an intermediate subgroup, then the intersections of all above subgroups with $H$ are Diophantine in $H$.
@@
By theorem 2 the following subgroups are Diophantine.
  • $\operatorname{C}^{\mathrm{us}}(G)(K)\,U_{\alpha}(K)$ for ultrashort $\alpha$ and $\Phi$ of type $\mathsf{BC}_{\ell}$;
  • $\operatorname{C}(G)(K)\,U_{2\mathrm{e}_{i}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,U_{2\mathrm{e}_{j}}(K)$ for $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ and $\Phi$ of type $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, $\ell\geq 2$;
  • $\mathrm{C}(G)(K)\,U_{\alpha}(K)$ for $\alpha$ long or $\alpha$ short and $\Phi$ of type $\mathsf{B}_{\ell}$, $\ell\geq 3$, $\mathsf{F}_{4}$, or $\mathsf{G}_{2}$.
Intersections of Diophantine subgroups, their products and their commutators with single elements are also Diophantine.
@@
$\operatorname{C}(G)(K)=\operatorname{C}(G)(K)\,U_{\alpha}(K)\cap\operatorname{C}(G)(K)\,U_{-\alpha}(K)$
@@
$\{g\in G(K)\mid{{}^{g}\!{t_{\alpha}(X)}}\subseteq\operatorname{C}(G)(K)\,U_{\alpha}(K)\}=G^{0}_{\{\beta\mid\angle(\alpha,\beta)\leq\pi/2\}}(K)$
@@
$U_{\alpha}(K)=\prod_{x\in X}\bigl[\operatorname{C}(G)(K)\,U_{\beta}(K),t_{\alpha-\beta}(x)\bigr]$
is Diophantine, where $\beta$ is long, $\angle(\alpha,\beta)=\frac{\pi}{3}$, and $X\subseteq\mathfrak{g}_{\alpha-\beta}$ is a finite generating set.
@@
$U_{\alpha}(K)=\operatorname{C}(G)(K)\,U_{\alpha}(K)\cap G_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}(G)(K)\,G_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr]$
is Diophantine, where $j\neq\pm i$ is another index and $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ is a finite generating set. It follows that all root subgroups are Diophantine for the type $\mathsf{F}_{4}$.
@@
$U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\prod_{x\in X,\,x^{\prime}\in X^{\prime}}\bigl[\bigl[\operatorname{C}(G)(K)\,U_{2\mathrm{e}_{j}}(K)\,U_{\mathrm{e}_{j}+\mathrm{e}_{k}}(K)\,U_{2\mathrm{e}_{k}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr],t_{\mathrm{e}_{j}-\mathrm{e}_{k}}(x^{\prime})\bigr],$
where $k\notin\{-i,i,-j,j\}$ is a new index and $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$, $X^{\prime}\subseteq\mathfrak{g}_{\mathrm{e}_{j}-\mathrm{e}_{k}}$ are finite generating sets. Next, the group
$U_{2\mathrm{e}_{i}}(K)=\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\cap U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}(G)\,U_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)]$
is Diophantine, where $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ is a finite generating set. If $\Phi$ is of type $\mathsf{BC}_{\ell}$, then
$U_{\mathrm{e}_{i}}(K)=\operatorname{C}^{\mathrm{us}}(G)(K)\,U_{\mathrm{e}_{i}}(K)\cap U_{\mathrm{e}_{i}-\mathrm{e}_{j}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}^{\mathrm{us}}(G)(K)\,U_{\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr]$
is also Diophantine, where $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ is a finite generating set.
Finally, consider the exceptional cases $\mathsf{C}_{2}$ and $\mathsf{BC}_{2}$. The group $\operatorname{C}(G)(K)\,U_{\mathrm{e}_{i}}(K)$ is Diophantine (in the case of $\mathsf{BC}_{2}$) because it contains $g\in\operatorname{C}^{\mathrm{us}}(K)\,U_{\mathrm{e}_{i}}(K)$ if and only if
$\bigl[g,t_{\mathrm{e}_{i}}(x)\bigr]\in\operatorname{C}(G)(K)\,U_{2\mathrm{e}_{i}}(K)$
for all $x\in X$, where $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\mathrm{e}_{i}}$ is a finite generating set. We have
$\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)=\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\,\prod_{x\in X}\bigl[\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}(K)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)\,U_{2\mathrm{e}_{j}}(K),t_{\mathrm{e}_{i}-\mathrm{e}_{j}}(x)\bigr]$
for a finite generating set $X\subseteq\mathfrak{g}_{\mathrm{e}_{i}-\mathrm{e}_{j}}$ by lemma 3, so $\operatorname{C}(G)\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}(K)$ is also Diophantine.
The last claim about intersections with $\mathrm{E}_{G}(K)\leq H\leq G(K)$ easily follows from explicit Diophantine formulas for $L(K)$, $\operatorname{C}(G)(K)$, $\operatorname{C}(G)(K)\,U_{\alpha}(K)$, and $U_{\alpha}(K)$. ∎
@@
Let $G$ be an isotropic reductive group scheme over $K$ with the root system $\Phi$ of rank at least $2$. Then the ring $K$ is e-interpretable in any intermediate subgroup $\mathrm{E}_{G}(K)\leq H\leq G(K)$.
@@
We assume that the type of $\Phi$ is neither $\mathsf{G}_{2}$ nor $\mathsf{BC}_{\ell}$ because these cases easily reduce to $\mathsf{A}_{2}$ and $\mathsf{C}_{\ell}$ respectively. Fix some finite $K$-module generating sets $X_{\alpha}\subseteq\mathfrak{g}_{\alpha}$ for all roots $\alpha$ and let $Y_{\alpha}=X_{\alpha}\cup\bigcup_{\beta,\alpha-\beta\in\Phi}[X_{\beta},X_{\alpha-\beta}]$.
Below we use that up to canonical isomorphisms the groups $\mathfrak{g}_{\alpha}$ are e-interpretable in $H$ by theorem 4 as Diophantine subgroups $U_{\alpha}(K)\leq H$ or as factor-groups $\operatorname{C}(H)\,U_{\alpha}(K)/\operatorname{C}(H)$. Under these e-interpretations all commutator maps $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ with $\alpha,\beta,\alpha+\beta\in\Phi$ have Diophantine graphs.
@@
$(k_{\alpha}\colon Y_{\alpha}\to\mathfrak{g}_{\alpha}(K))_{\alpha\in\Phi}$
@@
$k_{\alpha}(y_{\alpha}),y_{\beta}]=[y_{\alpha},k_{\beta}(y_{\beta})]\in\mathfrak{g}_{\alpha+\beta}$
@@
$k_{\alpha}(x_{\alpha}),x_{\beta}]=k_{\alpha+\beta}([x_{\alpha},x_{\beta}])\in\mathfrak{g}_{\alpha+\beta}$
@@
for $x_{\alpha}\in X_{\alpha}$, $x_{\beta}\in X_{\beta}$. Here $\alpha$ and $\beta$ are all roots such that $\alpha+\beta$ is also a root. Clearly, $\widetilde{K}$ is e-interpretable in $H$ as a set.
@@
Every $k\in K$ determines a corresponding element of $\widetilde{K}$ given by $y_{\alpha}\mapsto ky_{\alpha}$. Conversely, take $(k_{\alpha})_{\alpha\in\Phi}\in\widetilde{K}$. For every root $\alpha$ there is long $\beta\in\Phi$ such that $\alpha+\beta\in\Phi$, so $k_{\alpha}(y_{\alpha}),y_{\beta}]=[y_{\alpha},k_{\beta}(y_{\beta})]$. Recall that the Lie bracket $\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate (by lemma 3 if $\alpha$ is short). For any linear relation $\sum_{y_{\alpha}\in Y_{\alpha}}y_{\alpha}a_{\alpha}=0$ the equation implies that $\sum_{y_{\alpha}\in Y_{\alpha}}k_{\alpha}(y_{\alpha})a_{\alpha}=0$, so $k_{\alpha}$ continues to a unique linear map $k_{\alpha}\colon\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}$.
@@
The equations on $(k_{\alpha})_{\alpha}$ imply that $k_{\alpha}(x),y]=k_{\alpha+\beta}([x,y])=[x,k_{\beta}(y)]$. We claim that such maps $k_{\alpha}$ are necessarily scalar. It suffices to consider $\Phi$ of rank $2$ and to check the claim for only one root $\alpha$ (by lemma 3). This is clear if one of root subspaces has rank $1$. To prove the claim we can also assume that $G$ splits and choose convenient parametrizations of the root subspaces $\mathfrak{g}_{\alpha}$.
  • In the case $^{1}\!{\mathsf{A}_{3d-1,2}^{(d)}}$} we have
    $k_{\mathrm{e}_{1}-\mathrm{e}_{2}}(x)\,y=k_{\mathrm{e}_{1}-\mathrm{e}_{3}}(xy)=x\,k_{\mathrm{e}_{2}-\mathrm{e}_{3}}(y)$